each course at pease county community college is worth either 3 or 4 credit. The member of the women golf team are taken a total of 27 courses that are worth a total of 89 credit. How many 3 credit courses and how many 4 credit courses are being taken?
Answer by macston(5194)

.
T=number of 3 credit courses; F=number of 4 credit courses
T+F=27 Solve this for T.
T=27-F Use this to substitute for T.
3T+4F=89 Substitute for T from above.
3(27-F)+4F=89
81-3F+4F=89 Subtract 81 from each side.
F=8 ANSWER 11: There were eight - 4 credit courses.
T=27-F=27-8=19 ANSWER 2: There were 19 - 3 credit courses.
.
CHECK:
3T+4F=89
3(19)+4(8)=89
57+32=89
89=89